How long does it take for the energy stored in a capacitor in a series RC circuit to reach 75% of its maximum value? Express answer in terms of the time constant t = RC.
Answer #1

Energy stored on a capacitor U = (1/2) * C * V2

maximum energy U0   = 0.5 * C * V02

Hence 75% of maximum value U = 0.75 * U0   = 0.5 * C * V2

Dividing 0.75 =   V2 / V02

During charging V = V0 * (1 - e-t/τ) where τ (time constant) = RC

=> 0.75 = {V0 * (1 - e-t/τ) / V0}2

0.75 = (1 - e-t/τ)2

0.75 = 1 - e-t/τ

e-t/τ = 0.134

et/τ = 1/0.134

taking log t/R*C = ln (7.464)

time taken t = 2.01 * R * C